# Adaptive Accountability in Networked MAS:

### v1

@@ A.7 Proof of Corollary 2(Lorden Bound for Adaptive CUSUM)
Conditional on $h_{t}\!\to\!h^{\star}$ (Theorem 4) the adaptive rule behaves like a fixed‐threshold CUSUM with threshold $h^{\star}$. Page’s original result [page1954continuous] bounds the expected delay by $h^{\star}/(\Delta-\delta)$ when the post‐change mean is $\mu_{0}+\delta+\Delta$. Taking the supremum over $\tau^{\star}$ yields the worst‐case delay, completing the proof. $\Box$
@@ Appendix D. Mirror-Descent Alert Budget Allocator
where $z_{t}$ is the zero-one loss vector of fired alerts. Standard analysis [hazan2016introduction] yields regret $R_{T}\leq 2\sqrt{T\log|\Phi|}$; substituting $|\Phi|=5$ and $T=10^{6}$ gives $R_{T}\approx 10^{-2}$, hence the global alert budget $\bar{\alpha}=0.05$ is never exceeded by more than 0.01 in our runs.

### v2

@@ A.7 Proof of Corollary 2(Lorden Bound for Adaptive CUSUM)
Conditional on $h_{t}\!\to\!h^{\star}$ (Theorem 4) the adaptive rule behaves like a fixed‐threshold CUSUM with threshold $h^{\star}$. Page’s original result [46] bounds the expected delay by $h^{\star}/(\Delta-\delta)$ when the post‐change mean is $\mu_{0}+\delta+\Delta$. Taking the supremum over $\tau^{\star}$ yields the worst‐case delay, completing the proof. $\Box$
@@ Appendix D. Mirror-Descent Alert Budget Allocator
where $z_{t}$ is the zero-one loss vector of fired alerts. Standard analysis [23] yields regret $R_{T}\leq 2\sqrt{T\log|\Phi|}$; substituting $|\Phi|=5$ and $T=10^{6}$ gives $R_{T}\approx 10^{-2}$, hence the global alert budget $\bar{\alpha}=0.05$ is never exceeded by more than 0.01 in our runs.